### v1

@@ 4.3 Two-dimensional Helmholtz equation
where $\Omega=(=1,1)^{2}$ and $k=3$ is the unknown parameter, the exact solution is given by
@@ 4.4.1 Two-dimensional Problem
Figure 10: The sampling points at different levels for the two-dimensional interface problem.
@@
Figure 11: The numerical result of different levels for the two-dimensional interface problem.
The numerical results of different levels are given in Fig 11. Fig 11 displays a heat-map based on absolute error $|\bm{u}^{*}(\bm{x})-\bm{u}(\bm{x};\theta)|$. It can be observed that as the level increases, the approximation error decreases, demonstrating the effectiveness of the MLT method. After three levels of training, the relative errors of the final predicted solution are $e_{\infty}(\bm{u})=4.400\times 10^{-7}$ and $e_{2}(\bm{u})=8.197\times 10^{-8}$. In [4], the best relative $L^{2}$ error result for problem 4.4 in the two-dimensional case is $e_{2}(\bm{u})=8.769\times 10^{-7}$. In comparison, the accuracy of our method is superior.
@@ 4.4.2 Three-dimensional Problem
We take $d=3$ in Eq (40) and (41) here. In this experiment, the network structure was adopted with a size of 32×8. We sample 50000 points within $\Omega$ as the residual training set and 3000 points on $\partial\Omega$ as the boundary training set, while utilizing 160000 points for the test set, and we employ a three-level framework for training. In the first-level pre-training, we trained 20000 epochs using the SOAP method and 10000 epochs using the SSB method. Then, in the second-level training, we again trained 20000 epochs of the SOAP method and 10000 epochs of the SSB method. Finally, in the third-level training, we employed 20000 epochs of the SSB method. During training at the different levels, we employed the MLS method described in Section 3.1.1 and the MLT method described in Section 3.1.2.
@@
Figure 12: The sampling points at different levels for the three-dimensional interface problem.
@@
Figure 13: The numerical result of different levels for the three-dimensional interface problem. For clearer visualization, only 20000 random points within the computational domain were sampled when generating this plot.
@@
The numerical results of different levels are given in Fig 13. Fig 13 displays a heat-map based on absolute error $|\bm{u}^{*}(\bm{x})-\bm{u}(\bm{x};\theta)|$. It can be observed that as the level increases, the approximation error decreases, demonstrating the effectiveness of the MLT method. After three levels of training, the relative errors of the final predicted solution are $e_{\infty}(\bm{u})=8.707\times 10^{-7}$ and $e_{2}(\bm{u})=1.442\times 10^{-7}$. In [4], the best relative $L^{2}$ error result for problem 4.4 in the two-dimensional case is $e_{2}(\bm{u})=8.330\times 10^{-4}$. In comparison, the accuracy of our method is superior.
4.5 Two-dimensional Prandtl equation
@@
$\hskip-8.5359pt\begin{array}[]{r@{}l}\begin{aligned} u(0,y,t)&=u(1,y,t),\quad 0\leq y\leq 2,0\leq t\leq 0.16\\ v(0,x,y)&=v(1,y,t)\end{aligned}\end{array}$ (43)
@@
In this case, we will compare our numerical results with those reported in [8] (Fig 3.1). The governing equations and initial/boundary conditions used in our study are identical to those described in that reference. In Fig 3.1 of [8], the authors recast the inviscid Prandtl equations in conservation form 44a and explicitly discretize the resulting PDE in time.
@@
They evaluate the numerical x-flux for $u^{2}$ and adopt a Lax–Wendroff flux for the y-flux uv; v is computed from u by integrating 44b, discretized using central differences. Further details of the scheme can be found in [8].
@@
We sample 100000 points in $\Omega$ as the residual training set. And 100 points in each hyperplane on $\partial\Omega$ and 200000 points as the test set. Additionally, we uniformly sample 20000 points on the boundary to softly enforce the periodic boundary condition in x, and another 10000 points to softly enforce the initial condition. At the same time, the boundary condition for $v(x,y,t)$ at y = 0 is imposed exactly (hard enforcement). In this experiment, during the first-level pre-training, we trained 20000 epochs using the SOAP method. In the second-level training, we trained 20000 epochs of the SOAP method and 30000 epochs of the SSB method. Finally, in the third-level training, we employed 50000 epochs of the SSB method. The distributions of residual sampling points as well as the behavior of the loss function across different training levels, are summarized in Fig 14.
@@
(d) Loss
Figure 14: (a)-(c) the sampling points at different levels for two-dimensional Prandtl equation; (d) the behavior of the loss function across different training levels.
In Fig 15 , we compare the numerical results derived from GLP sampling with those from uniform random sampling after 7500 epochs of Adam training and 10000 epochs of LBFGS training. This comparison aims to demonstrate the efficiency of our method for the same number of training points. In Fig 16, we present surface plots of the neural network predictions for $u(x,y,t)$ and $v(x,y,t)$ at t = 0.16 at different training levels. In Figure 12, we show two-dimensional graphs of the predictions of $u(x,y,t)$ and $v(x,y,t)$ at y = 1 for $t=0,0.08,0.16$, again at each level. By comparing Fig 3.1 in [8], it is immediately apparent that the results obtained with our method are are comparable to the high-precision numerical scheme employed in [8]. A closer inspection of Fig 15 further reveals that our approach captures the spike of $v(x,y,t)$ even more accurately.
@@
(a) $uX1T$ at 1st level
(b) $uX1T$ at 2nd level
(c) $uX1T$ at 3rd level
(d) $vX1T$ at 1st level
(e) $vX1T$ at 2nd level
(f) $vX1T$ at 3rd level
Figure 16: The plots of $uX1T$ and $vX1T$ at different levels for two-dimensional Prandtl equation.

### v2

@@ 4.4 Two-dimensional Helmholtz equation
where $\Omega=(=1,1)^{2}$ and $k=3$, the exact solution is given by
@@ 4.5.1 Two-dimensional Problem
Figure 13: The sampling points at different levels for the two-dimensional Poisson equation with the solution Eq (50).
@@
The numerical results of different levels are given in Fig 14. Fig 14 displays a heat-map based on absolute error $|\bm{u}^{*}(\bm{x})-\bm{u}(\bm{x};\theta)|$. It can be observed that as the level increases, the approximation error decreases, demonstrating the effectiveness of the MLT method. After three levels of training, the relative errors of the final predicted solution are $e_{\infty}(\bm{u})=4.400\times 10^{-7}$ and $e_{2}(\bm{u})=8.197\times 10^{-8}$. In [4], the best relative $L^{2}$ error result for problem 4.5 in the two-dimensional case is $e_{2}(\bm{u})=8.769\times 10^{-7}$. In comparison, the accuracy of our method is superior.
@@ 4.5.2 Three-dimensional Problem
We take $d=3$ in Eq (49) and (50) here. In this experiment, the network structure was adopted with a size of 32×8. We sample 50000 points within $\Omega$ as the residual training set and 3000 points on $\partial\Omega$ as the boundary training set, while utilizing 160000 points for the test set, and we employ a three-level framework for training. In the first-level pre-training, we trained 20000 epochs using the SOAP method and 10000 epochs using the SSB method. Then, in the second-level training, we again trained 20000 epochs of the SOAP method and 10000 epochs of the SSB method. Finally, in the third-level training, we employed 20000 epochs of the SSB method.
During training at the different levels, we employed the MLS method described in Section 3.1.1. Fig 15 illustrates the effect of the MLS method across the various levels; it can be seen that as the level increases, the points progressively concentrate toward the origin, because the solution has very low regularity in the vicinity of the origin. It can be seen that the MLS method is able to capture this property effectively.
@@
Figure 15: The sampling points at different levels for the three-dimensional Poisson equation with the solution Eq (50).
@@
Figure 16: The numerical result of different levels for the three-dimensional Poisson equation with the solution Eq (50). For clearer visualization, only 20000 random points within the computational domain were sampled when generating this plot.
@@
The numerical results of different levels are given in Fig 16. Fig 16 displays a heat-map based on absolute error $|\bm{u}^{*}(\bm{x})-\bm{u}(\bm{x};\theta)|$. It can be observed that as the level increases, the approximation error decreases, demonstrating the effectiveness of the MLT method. After three levels of training, the relative errors of the final predicted solution are $e_{\infty}(\bm{u})=8.707\times 10^{-7}$ and $e_{2}(\bm{u})=1.442\times 10^{-7}$. In [4], the best relative $L^{2}$ error result for problem 4.5 in the three-dimensional case is $e_{2}(\bm{u})=8.330\times 10^{-4}$. In comparison, the accuracy of our method is superior.
4.6 Two-dimensional Poisson Equation with L shaped domain
For the following Poisson equation in two-dimension
[…]
where the domain is L shaped $\Omega=(-1,1)^{2}\backslash(0,1)\times(-1,0)$, the exact solution is chosen as
$\hskip-8.5359pt\begin{array}[]{r@{}l}\begin{aligned} u=r^{2/3}sin\frac{2\theta}{3}-\frac{x^{2}+y^{2}-2}{4}.\end{aligned}\end{array}$ (52)
where r and $\theta$ are the coordinates in the polar coordinate system, i.e., $r=\sqrt{x^{2}+y^{2}},\tan(\theta)=\frac{y}{x}$.
In this experiment, we sample $7500$ points within $\Omega$ as the residual training set and 1000 points on $\partial\Omega$ as the boundary training set, while utilizing 119600 points for the test set, and we employ a three-level framework for training . In the first-level pre-training, we trained 20000 epochs using the SOAP method. Then, in the second-level training, we again trained 10000 epochs of the SOAP method and 5000 epochs of the SSB method. Finally, in the third-level training, we employed 2500 epochs of the SOAP method and 12500 epochs of the SSB method.
During training at the different levels, we employed the MLS method described in Section 3.1.1. The large gradient of the solution near the origin is a major cause of difficulty in solving the problem. To mitigate the impact of the gradient term, we adjusted the parameters ($\gamma=0.001$,$c=0.1$) in the MLS method. Fig 17 illustrates the effect of the MLS method across the various levels; it can be seen that as the level increases, the points progressively concentrate toward $(0,0)$, because the solution has very low regularity in the vicinity of $(0,0)$. It can be seen that the MLS method is able to capture this property effectively.
(a) 1st level
(b) 2nd level
(c) 3rd level
Figure 17: The sampling points at different levels for the two-dimensional Poisson equation with the solution Eq (52).
The numerical results of different levels are given in Fig 18. Fig 18 illustrates figures based on the predicted solutions $\bm{u}(\bm{x};\theta)|$ and the absolute error $|\bm{u}^{*}(\bm{x})-\bm{u}(\bm{x};\theta)|$. It can be observed that as the level increases, the approximation error decreases, demonstrating the effectiveness of our multi-level deep framework.
(a) 1st prediction
(b) 2nd prediction
(c) 3rd prediction
(d) 1st error
(e) 2nd error
(f) 3rd error
Figure 18: The numerical result of different levels for the two-dimensional Poisson equation with the solution Eq (52). (a)–(c) The predicted $\bm{u}(\bm{x};\theta)$. (d)–(f) Performance of the absolute error $|\bm{u}^{*}(\bm{x})-\bm{u}(\bm{x};\theta)|$.
In Fig 19, we compare the commonly used RAD method in PINNs under the same number of training epochs. Specifically, the training process consists of pre-training with 20000 iterations of Adam optimizer, followed by sampling with the RAD method, finally trains again for 12500 Adam epochs and 17500 L-BFGS epochs. Fig 19 illustrates the variation of the relative errors during the training process, from which the superiority of our method can be observed. Ultimately, the relative errors of our method in solving the two-dimensional Poisson equation with the solution Eq (52) are $e_{\infty}(\bm{u})=2.209\times 10^{-3}$ and $e_{2}(\bm{u})=2.716\times 10^{-5}$.
(a) performance of $e_{\infty}(u)$
(b) performance of $e_{2}(u)$
4.7 Fourth-order Biharmonic equation
For the following Biharmonic equation in two-dimension
[…]
where the domain is the unit disk $\Omega=\left\{(x,y)|x^{2}+y^{2}<1\right\}$, and the source term is the Dirac function $\delta_{(0,0)}$. The exact solution is given by
$\hskip-8.5359pt\begin{array}[]{r@{}l}\begin{aligned} u=\frac{r^{2}}{8\pi}\ln r+c_{1}r^{2}+c_{2},\ c_{1}=\frac{1}{4+2\epsilon_{2}}(-\frac{1}{2\pi}-\frac{\epsilon_{2}}{8\pi}),\ c_{2}=-c_{1}+\frac{\epsilon_{1}}{2\pi}.\end{aligned}\end{array}$ (54)
where r and $\theta$ are the coordinates in the polar coordinate system.
In this experiment, we take $\epsilon_{1}=\epsilon_{2}=10^{-5}$ and sample $100\times 100$ points within $\Omega$ as the residual training set and 256 points on $\partial\Omega$ as the boundary training set, while utilizing $400\times 400$ points for the test set, and we employ a three-level framework for training . In the first-level pre-training, we trained 20000 epochs using the SOAP method. Then, in the second-level training, we again trained 5000 epochs of the SOAP method and 10000 epochs of the SSB method. Finally, in the third-level training, we employed 300 epochs the SSB method.
During training at the different levels, we employed the MLS method described in Section 3.1.1. To enhance the concentration trend of the sampling points, we adjusted the parameters ($\gamma=0$, $s=1$) in the MLS method. Fig 20 illustrates the effect of the MLS method across the various levels; it can be seen that as the level increases, the points progressively concentrate toward $(0,0)$, because the solution has very low regularity in the vicinity of $(0,0)$. It can be seen that the MLS method is able to capture this property effectively.
(a) 1st level
(b) 2nd level
(c) 3rd level
Figure 20: The sampling points at different levels for the two-dimensional Biharmonic equation with the solution Eq (54).
The numerical results of different levels are given in Fig 21. Fig 21 illustrates figures based on the predicted solutions $\bm{u}(\bm{x};\theta)|$ and the absolute error $|\bm{u}^{*}(\bm{x})-\bm{u}(\bm{x};\theta)|$. It can be observed that as the level increases, the approximation error decreases, demonstrating the effectiveness of our multi-level deep framework.
(a) 1st prediction
(b) 2nd prediction
(c) 3rd prediction
(d) 1st error
(e) 2nd error
(f) 3rd error
Figure 21: The numerical result of different levels for the two-dimensional Biharmonic equation with the solution Eq (54). (a)–(c) The predicted $\bm{u}(\bm{x};\theta)$. (d)–(f) Performance of the absolute error $|\bm{u}^{*}(\bm{x})-\bm{u}(\bm{x};\theta)|$.
Ultimately, the relative errors of our method in solving the two-dimensional Biharmonic equation with the solution Eq (54) are $e_{\infty}(\bm{u})=1.651\times 10^{-4}$ and $e_{2}(\bm{u})=4.774\times 10^{-5}$.
4.8 Two-dimensional inviscid Prandtl equation
@@
$\hskip-8.5359pt\begin{array}[]{r@{}l}\begin{aligned} u(0,y,t)&=u(1,y,t),\quad 0\leq y\leq 2,0\leq t\leq 0.16\\ v(0,y,t)&=v(1,y,t)\end{aligned}\end{array}$ (56)
@@
In this case, we will compare our numerical results with those reported in [8] (Fig 3.1). The governing equations and initial/boundary conditions used in our study are identical to those described in that reference. In Fig 3.1 of [8], the authors recast the inviscid Prandtl equations in conservation form 57a and explicitly discretize the resulting PDE in time.
@@
They evaluate the numerical x-flux for $u^{2}$ and adopt a Lax–Wendroff flux for the y-flux uv; v is computed from u by integrating 57b, discretized using central differences. Further details of the scheme can be found in [8].
@@
We sample 100000 points in $\Omega$ as the residual training set. And 100 points in each hyperplane on $\partial\Omega$. Additionally, we uniformly sample 20000 points on the boundary to softly enforce the periodic boundary condition in x, and another 10000 points to softly enforce the initial condition. At the same time, the boundary condition for $v(x,y,t)$ at y = 0 is imposed exactly (hard enforcement). In this experiment, during the first-level pre-training, we trained 20000 epochs using the SOAP method. In the second-level training, we trained 20000 epochs of the SOAP method and 30000 epochs of the SSB method. Finally, in the third-level training, we employed 50000 epochs of the SSB method. The distributions of residual sampling points are summarized in Fig 22.
@@
Figure 22: (a)-(c) The sampling points at different levels for two-dimensional inviscid Prandtl equation
In Fig 23 , we present surface plots of the neural network predictions for $u(x,y,t)$ and $v(x,y,t)$ at t = 0.16 at different training levels. In Figure 24, we show two-dimensional graphs of the predictions of $u(x,y,t)$ and $v(x,y,t)$ at y = 1 for $t=0,0.08,0.16$, again at each level. By comparing Fig 3.1 in [8], it is immediately apparent that the results obtained with our method are are comparable to the high-precision numerical scheme employed in [8]. A closer inspection of Fig 23 further reveals that our approach captures the spike of $v(x,y,t)$ even more accurately.
(a) $u(x,y,0.16)$ at 1st level
(b) $u(x,y,0.16)$ at 2nd level
(c) $u(x,y,0.16)$ at 3rd level
(d) $v(x,y,0.16)$ at 1st level
(e) $v(x,y,0.16)$ at 2nd level
(f) $v(x,y,0.16)$ at 3rd level
Figure 23: The surface plots of $u(x,y,0.16)$ and $v(x,y,0.16)$ at different levels for two-dimensional inviscid Prandtl equation.
(a) $u(x,1,t)$ at 1st level
(b) $u(x,1,t)$ at 2nd level
(c) $u(x,1,t)$ at 3rd level
(d) $v(x,1,t)$ at 1st level
(e) $v(x,1,t)$ at 2nd level
(f) $v(x,1,t)$ at 3rd level
Figure 24: The plots of $u(x,1,t)$ and $v(x,1,t)$ at different levels for two-dimensional inviscid Prandtl equation.
4.9 Two-dimensional viscous Prandtl equation
For the following unsteady Prandtl equations with zero pressure,
[…]
where $\nu=0.1$.
In this viscous problem, as [8], we specify a finite computational domain $\Omega=[0,1]\times[0,2]\times[0,0.16]$.
In this case, we will compare our numerical results with those reported in [8] (Fig 4.3). The governing equations and initial/boundary conditions used in our study are identical to those described in that reference. Further details of the scheme can be found in [8].
We sample 100000 points in $\Omega$ as the residual training set. Additionally, we uniformly sample 40000 points on the boundary to softly enforce the boundary condition in x, and another 5000 points to softly enforce the initial condition. At the same time, the boundary condition for $v(x,y,t)$ at y = 0 is imposed exactly (hard enforcement). In this experiment, during the first-level pre-training, we trained 20000 epochs using the SOAP method. In the second-level training, we trained 20000 epochs of the SOAP method and 30000 epochs of the SSB method. Finally, in the third-level training, we employed 50000 epochs of the SSB method. The distributions of residual sampling points as well as the behavior of the loss function across different training levels, are summarized in Fig 25.
(a) 1st level
(b) 2nd level
(c) 3rd level
Figure 25: (a)-(c) the sampling points at different levels for two-dimensional viscous Prandtl equation.
In Fig 26, we present surface plots of the neural network predictions for $u(x,y,t)$ and $v(x,y,t)$ at t = 0.16 at different training levels. In Figure 27 , we show two-dimensional graphs of the predictions of $u(x,y,t)$ and $v(x,y,t)$ at x = 0.5 for $t=0.08,0.16$ at each level. In Figure 28 , we show two-dimensional graphs of the predictions of $u(x,y,t)$ and $v(x,y,t)$ at y = 1 for $t=0,0.08,0.16$, again at each level.
By comparing Fig 4.3 in [8], it is immediately apparent that the results obtained with our method are are comparable to the high-precision numerical scheme employed in [8]. A closer inspection of Fig 26 further reveals that our approach captures the spike of $v(x,y,t)$ even more accurately.
@@ 4.9 Two-dimensional viscous Prandtl equation
(a) $\bm{u}(0.5,y,t)$ at 1st level
(b) $\bm{u}(0.5,y,t)$ at 2nd level
(c) $\bm{u}(0.5,y,t)$ at 3rd level
(d) $\bm{v}(0.5,y,t)$ at 1st level
(e) $\bm{v}(0.5,y,t)$ at 2nd level
(f) $\bm{v}(0.5,y,t)$ at 3rd level
Figure 27: The plots of $\bm{u}(0.5,Y,t)$ and $\bm{v}(0.5,Y,t)$ at different levels for two-dimensional viscous Prandtl equation.
(a) $\bm{u}(x,1,t)$ at 1st level
(b) $\bm{u}(x,1,t)$ at 2nd level
(c) $\bm{u}(x,1,t)$ at 3rd level
(d) $\bm{v}(x,1,t)$ at 1st level
(e) $\bm{v}(x,1,t)$ at 2nd level
(f) $\bm{v}(x,1,t)$ at 3rd level
Figure 28: The plots of $\bm{u}(x,1,t)$ and $\bm{v}(x,1,t)$ at different levels for two-dimensional viscous Prandtl equation.